OpenStudy (lena772):

Madeline was completing the square and her work is shown below. Identify the line where she made her mistake. f(x) = 3x^2 + 6x − 7 Line 1: f(x) = 3(x^2 + 2x) − 7 Line 2: f(x) = 3(x^2 + 2x + 1) − 7 − 1 Line 3: f(x) = 3(x + 1)^2 − 8 Line 1 Line 2 Line 3 She did not make any mistakes

ganeshie8 (ganeshie8):

whats ur guess

OpenStudy (lena772):

LiNE 3

ganeshie8 (ganeshie8):

good guess, but no

OpenStudy (lena772):

ok

ganeshie8 (ganeshie8):

Clearly there is nothing wrong in the Line1, she is just factoring out 3 form first two terms

ganeshie8 (ganeshie8):

Line 1: f(x) = 3(x^2 + 2x) − 7 Line 2: f(x) = 3(x^2 + 2x + 1) − 7 − 1

OpenStudy (lena772):

Line 2 then?

ganeshie8 (ganeshie8):

yes lol, if its not line1 and line3, then it has to be line2 u dont wana knw whats wrong wid line2 ?

OpenStudy (lena772):

Is it the -1 at the end?

ganeshie8 (ganeshie8):

yes, it has to be -3 instead

OpenStudy (lena772):

thanks

ganeshie8 (ganeshie8):

yw